OpenStudy (ajanijones):

I GIVE MEDALS 30% of a number is 120. What is 75% of the number? A. 300 B. 90 C. 27 D. 48

OpenStudy (anonymous):

75% of 120, is 90

OpenStudy (anonymous):

SO it's B

OpenStudy (ajanijones):

Thanks. you realy are smart very smart lol i cant think today

OpenStudy (anonymous):

xD

OpenStudy (ranga):

Let x be the number. 0.3x = 120 x = 120/0.3 75% of x is: .75x = .75 * 120 / 0.3 = ?

OpenStudy (campbell_st):

find 5% divide both by 6 5% = 20 to find 75% multiply both by 15 5 x 15 = 75% 15 x 20 = Answer

OpenStudy (campbell_st):

p.s.... 90 isn't the answer

OpenStudy (anonymous):

.3 x = 120 x = 400 3/4 400 = 300

OpenStudy (anonymous):

*raises hand* where did the 3 come from......?

OpenStudy (anonymous):

30% = 0.3

OpenStudy (anonymous):

oh i didnt see the point xD